# Balancing Task-invariant Interaction and Task-specific Adaptation for Unified Image Fusion

### v1

@@ Abstract
Unified image fusion aims to integrate complementary information from multi-source images, enhancing image quality through a unified framework applicable to diverse fusion tasks. While treating all fusion tasks as a unified problem facilitates task-invariant knowledge sharing, it often overlooks task-specific characteristics, thereby limiting the overall performance. Existing general image fusion methods incorporate explicit task identification to enable adaptation to different fusion tasks. However, this dependence during inference restricts the model’s generalization to unseen fusion tasks. To address these issues, we propose a novel unified image fusion framework named “TITA”, which dynamically balances both Task-invariant Interaction and Task-specific Adaptation. For task-invariant interaction, we introduce the Interaction-enhanced Pixel Attention (IPA) module to enhance pixel-wise interactions for better multi-source complementary information extraction. For task-specific adaptation, the Operation-based Adaptive Fusion (OAF) module dynamically adjusts operation weights based on task properties. Additionally, we incorporate the Fast Adaptive Multitask Optimization (FAMO) strategy to mitigate the impact of gradient conflicts across tasks during joint training. Extensive experiments demonstrate that TITA not only achieves competitive performance compared to specialized methods across three image fusion scenarios but also exhibits strong generalization to unseen fusion tasks.
@@ 1 Introduction
Image fusion encompasses diverse scenarios, such as infrared-visible image fusion (IVF), multi-exposure image fusion (MEF), and multi-focus image fusion (MFF). These fusion tasks exhibit distinct characteristics. For instance, IVF emphasizes salient thermal features from infrared images while preserving texture details from visible images. MEF balances color and brightness across over-exposed and under-exposed images, whereas MFF focuses on extracting sharp regions from near-focus and far-focus images. Despite specific objectives for feature extraction may vary across tasks, all image fusion methods share a fundamental goal: integrating complementary information from source images to generate a high-quality fused image.
@@
  • An Interaction-enhanced Pixel Attention (IPA) block that discovers multi-source feature correlations is introduced to explore the task-invariant properties. An Operation-based Adaptive Fusion (OAF) module that dynamically adjusts fusion operations is designed to explore the task-specific properties without explicit task identification. The framework is reinforced by Fast Adaptive Multitask Optimization (FAMO) strategy to mitigate the impact of objective conflicts.
@@ 3.2 Task-Specific Feature Fusion
Specifically, the OAF module consists of three parallel operation branches: (i) HPF branch: applies spatially-variant high-pass filtering to capture high-frequency details, enhancing texture and edge information. (ii) ADD branch: performs residual addition for overall information enhancement. (iii) MUL branch: performs element-wise multiplication to facilitate nonlinear feature interactions to capture complex feature relationships. The dynamic weights of these branches reflect the physical characteristics of different fusion tasks (detailed in Sec. 4.3.3). Instead of manually assigning fixed weights to each fusion task, we employ a dynamic weight prediction network that takes task-specific features as input and predicts adaptive weights to determine the optimal weight distribution. As illustrated in Fig. 1(c), given multi-source features $X_{1},X_{2}$, the hypernetworks generate operands $P_{h}=\phi_{\theta_{h}}(X),P_{a}=\phi_{\theta_{a}}(X),P_{m}=\phi_{\theta_{m}}(X)$. And the resultant features $\hat{X}_{h}=X*P_{h},\hat{X}_{a}=X+P_{a},\hat{X}_{m}=X\odot P_{m}$ are then dynamically weighted by the weights $W=\phi_{\theta_{w}}(X_{1},X_{2})$. Then the output feature $X_{f}$ can be obtained as:
@@
Figure 2: The gradient conflict angles and the gradient norms on three fusion tasks when TITA operates without multi-objective optimization. We observed gradient conflicts occurring frequently during training, and the case shown here (iteration=$3000$) was selected as a representative example.
@@ 4.1 Experimental Setup
We use an Adam optimizer with learning rate set to $2e^{-5}$ for model parameter optimization, and an Adam optimizer with learning rate set to $0.025$ and weight decay set to $1e^{-3}$ for weighting logits optimization (introduced by FAMO). The batch size is set to $8$ and the number of iterations is set to $20000$. Training data is unevenly distributed across fusion tasks due to varying acquisition costs and difficulties. To address this imbalance, we uniformly sample data from each task for each iteration. The proposed method is implemented by Pytorch and all experiments are conducted on a device equipped with a NVIDIA RTX 3090 GPU.
@@
We follow TC-MoA [62] to construct the training and testing datasets. The training dataset consists of $12025$ infrared and visible image pairs from LLVIP dataset [12], $589$ multi-exposure image pairs from SCIE dataset [3], $710$ multi-focus image pairs from RealMFF dataset [55], and $90$ multi-focus image pairs from MFI-WHU dataset [53]. Although the trainig data size varies across different fusion tasks, all tasks are sampled equally to mitigate the imbalance. For evaluation on IVF, the testing dataset includes $70$ infrared and visible image pairs from LLVIP dataset [12]. For evaluation on MEF, the testing dataset includes $100$ multi-exposure image pairs from MEFB dataset [56]. And for evaluation on MFF, the testing dataset contains $20$ multi-focus image pairs from Lytro dataset [32], $13$ multi-focus image pairs from MFFW dataset [46], and $30$ multi-focus image pairs from MFI-WHU dataset [53]. Besides, we validate the generalization of the proposed unified image fusion method on Harvard dataset  [^1] http://www.med.harvard.edu/AANLIB/home.html for medical image fusion, and Quickbird dataset  [^2] https://www.satimagingcorp.com/satellite-sensors/quickbird/ for pan-sharpening following PMGI [51].
@@ 4.2.4 Generalization on Other Fusion Tasks
We test TITA on two unseen fusion tasks, that is, medical image fusion (MIF) and pan-sharpening (PAN), and compare with six unified image fusion methods. Note that for PMGI [51], these two tasks are known. Fig. 6 shows that TITA demonstrates strong adaptability to different types of fusion scenarios, benefiting from its task-agnostic design. In contrast, FusionDN and CCF fail to generalize and collapse when encountering unseen tasks. This further verifies that the common goal of pursuing the maximization of the amount of information among different fusion tasks enables the unified method to be effectively generalized to various scenarios.
Table 4: The ablation study on three main components of the proposed method.
@@ 4.3.2 Analysis on Task-invariant Integration
This paragraph mainly focuses on the verification of the proposed IPA module and the ISF module. First, we compare the proposed Interaction-enhanced (IeSF) module which replaces all the intra-domain fusion with the inter-domain fusion, with the baseline-ST (with original SwinFusion (SF) module) and the Interaction-reduced (IrSF) module which replaces all the inter-domain fusion with the intra-domain fusion. The key difference among these modules lies in the balance between cross-attention and self-attention operations. Specifically, the number of cross-attention mechanisms follows the order: IeSF $>$ SF $>$ IrSF, highlighting the increasing level of multi-source interactions. The results shown in Tab. 5 indicate that increasing the amount of cross-attention operations leads to improved performance. Then we compare IPA with PA [11] and Token Exchange (TE) [39], which is introduced in the Appendix A.3. The results shown in Tab. 5 demonstrate the effectiveness of the proposed IPA module.
@@ 4.3.3 Analysis on Task-specific Adaptation
To illustrate the importance of each operation branch and the dynamic weight prediction (DW) in the OAF module, we conduct an ablation study, as shown in Tab. 6. The results indicate that both the DW strategy and the three operation branches contribute significantly to the overall performance. Among these branches, the MUL branch proves to be the most crucial, which is reasonable given that image fusion inherently involves many non-linear operations. This highlights the necessity of incorporating diverse operations to enhance the adaptive feature fusion. The visualization of the dynamic weights is provided in Appendix B.2.
@@ 5 Conclusion
In this work, we propose a unified image fusion framework that effectively integrates both task-invariant and task-specific properties to enhance fusion performance and generalization across different fusion tasks. Future work will explore further improvements through dynamic expansion to other unseen fusion scenarios and integrate advanced learning techniques to develop stronger cross-modal representations, enhancing fusion quality and adaptability across diverse applications.
@@ Appendix A Supplementary Details
A.1 Optimization Objectives
@@
A.2 Additional Details of FAMO
@@ A.3 Token Exchange
where $\mathbb{I}$ is an indicator, and the threshold $\gamma$ is set to $0.02$ according to the paper.
@@ Appendix B More Results and Analysis
Table 7: The ablation study on MEF task.
@@
Table 8: The ablation study on MFF task.
@@
B.1 Ablation study on MEF and MFF tasks
@@
B.2 Visualization of TA
The weights for each branch of OAF on all fusion tasks is visualized in Tab. 9. It can be found that the weight of the HPF branch is small because high-frequency details constitute only a small portion of the overall information. However, as shown in the ablation study on task-invariant integration, the high-frequency details play a crucial role in enhancing performance.

### v2

@@ Abstract
Unified image fusion aims to integrate complementary information from multi-source images, enhancing image quality through a unified framework applicable to diverse fusion tasks. While treating all fusion tasks as a unified problem facilitates task-invariant knowledge sharing, it often overlooks task-specific characteristics, thereby limiting the overall performance. Existing general image fusion methods incorporate explicit task identification to enable adaptation to different fusion tasks. However, this dependence during inference restricts the model’s generalization to unseen fusion tasks. To address these issues, we propose a novel unified image fusion framework named “TITA”, which dynamically balances both Task-invariant Interaction and Task-specific Adaptation. For task-invariant interaction, we introduce the Interaction-enhanced Pixel Attention (IPA) module to enhance pixel-wise interactions for better multi-source complementary information extraction. For task-specific adaptation, the Operation-based Adaptive Fusion (OAF) module dynamically adjusts operation weights based on task properties. Additionally, we incorporate the Fast Adaptive Multitask Optimization (FAMO) strategy to mitigate the impact of gradient conflicts across tasks during joint training. Extensive experiments demonstrate that TITA not only achieves competitive performance compared to specialized methods across three image fusion scenarios but also exhibits strong generalization to unseen fusion tasks. The source codes are released at https://github.com/huxingyuabc/TITA.
@@ 1 Introduction
Image fusion encompasses diverse scenarios, such as infrared-visible image fusion (IVF), multi-exposure image fusion (MEF), and multi-focus image fusion (MFF). These fusion tasks exhibit distinct characteristics. For instance, IVF emphasizes salient thermal features from infrared images while preserving texture details from visible images. MEF balances color and brightness across over- and under-exposed images, whereas MFF focuses on extracting sharp regions from near- and far-focus images. Despite variations in feature extraction objectives, all image fusion methods share a fundamental goal: integrating complementary information and generating high-quality fused images.
@@
  • An Interaction-enhanced Pixel Attention (IPA) block that discovers multi-source feature correlations is introduced to explore task-invariant properties. An Operation-based Adaptive Fusion (OAF) module that dynamically adjusts fusion operations is designed to capture task-specific properties without explicit task identification. The framework is reinforced by Fast Adaptive Multitask Optimization (FAMO) strategy to mitigate objective conflicts.
@@ 3.2 Task-Specific Feature Fusion
Specifically, the OAF module consists of three parallel operation branches: (i) HPF branch: applies spatially-variant high-pass filtering to capture high-frequency details, enhancing texture and edge information. (ii) ADD branch: performs residual addition for overall information enhancement. (iii) MUL branch: performs element-wise multiplication to facilitate nonlinear feature interactions to capture complex feature relationships. (For details, please refer Appendix A.1.) The dynamic weights of these branches reflect the physical characteristics of different fusion tasks. Instead of manually assigning fixed weights to each fusion task, we employ a dynamic weight prediction network that takes task-specific features as input and predicts adaptive weights to determine the optimal weight distribution. As illustrated in Fig. 1(c), given multi-source features $X_{1},X_{2}$, the hypernetworks generate operands $P_{h}=\phi_{\theta_{h}}(X),P_{a}=\phi_{\theta_{a}}(X),P_{m}=\phi_{\theta_{m}}(X)$. And the resultant features $\hat{X}_{h}=X*P_{h},\hat{X}_{a}=X+P_{a},\hat{X}_{m}=X\odot P_{m}$ are then dynamically weighted by the weights $W=\phi_{\theta_{w}}(X_{1},X_{2})$. Then the output feature $X_{f}$ can be obtained as:
@@
Figure 2: The gradient conflict angles and the gradient norms on three fusion tasks when TITA operates without multi-objective optimization. We observed gradient conflicts occurring frequently during training, and the case shown here (iteration=300030003000) was selected as a representative example.
@@ 4.1 Experimental Setup
We use an Adam optimizer with learning rate set to $2e^{-5}$ for model parameter optimization, and an Adam optimizer with learning rate set to 0.0250.0250.025 and weight decay set to $1e^{-3}$ for weighting logits optimization (introduced by FAMO). The batch size is set to 888 and the number of iterations is set to 200002000020000. Training data is unevenly distributed across fusion tasks due to varying acquisition costs and difficulties. To address this imbalance, we uniformly sample data from each task for each iteration. The proposed method is implemented by Pytorch and all experiments are conducted on a device equipped with a NVIDIA RTX 3090 GPU.
@@
We follow TC-MoA [62] to construct the training and testing datasets. The training dataset consists of 120251202512025 infrared and visible image pairs from LLVIP dataset [12], 589589589 multi-exposure image pairs from SCIE dataset [3], 710710710 multi-focus image pairs from RealMFF dataset [55], and 909090 multi-focus image pairs from MFI-WHU dataset [53]. Although the trainig data size varies across different fusion tasks, all tasks are sampled equally to mitigate the imbalance. For evaluation on IVF, the testing dataset includes 707070 infrared and visible image pairs from LLVIP dataset [12]. For evaluation on MEF, the testing dataset includes 100100100 multi-exposure image pairs from MEFB dataset [56]. And for evaluation on MFF, the testing dataset contains 202020 multi-focus image pairs from Lytro dataset [32], 131313 multi-focus image pairs from MFFW dataset [46], and 303030 multi-focus image pairs from MFI-WHU dataset [53]. Besides, we validate the generalization of the proposed unified image fusion method on Harvard dataset  [^1] http://www.med.harvard.edu/AANLIB/home.html for medical image fusion, and Quickbird dataset  [^2] https://www.satimagingcorp.com/satellite-sensors/quickbird/ for pan-sharpening following PMGI [51].
@@ 4.2.4 Generalization on Other Fusion Tasks
We test TITA on two unseen fusion tasks, that is, medical image fusion (MIF) and pan-sharpening (PAN), and compare with six unified image fusion methods. Note that for PMGI [51], these two tasks are known. Fig. 6 shows that TITA demonstrates strong adaptability to different types of fusion scenarios, benefiting from its task-agnostic design. In contrast, FusionDN and CCF fail to generalize and collapse when encountering unseen tasks. This further verifies that the common goal of maximizing the amount of information among different fusion tasks enables the unified method to be effectively generalized to various scenarios.
Table 4: The ablation study on three main components.
@@ 4.3.2 Analysis on Task-invariant Integration
This paragraph mainly focuses on the verification of the proposed IPA module and the ISF module. First, we compare the proposed Interaction-enhanced (IeSF) module which replaces all the intra-domain fusion with the inter-domain fusion, with the baseline-ST (with original SwinFusion (SF) module) and the Interaction-reduced (IrSF) module which replaces all the inter-domain fusion with the intra-domain fusion. The key difference among these modules lies in the balance between cross-attention and self-attention operations. Specifically, the number of cross-attention mechanisms follows the order: IeSF $>$ SF $>$ IrSF, highlighting the increasing level of multi-source interactions. The results shown in Tab. 5 indicate that increasing the amount of cross-attention operations leads to improved performance. (Results on MEF and MFF tasks in Appendix B.2 also validate this.) Then we compare IPA with PA [11] and Token Exchange (TE) [39], which is introduced in the Appendix A.4. The results shown in Tab. 5 demonstrate the effectiveness of the proposed IPA module.
@@ 4.3.3 Analysis on Task-specific Adaptation
To illustrate the importance of each operation branch and the dynamic weight prediction (DW) in the OAF module, we conduct an ablation study, as shown in Tab. 6. The results indicate that both the DW strategy and the three operation branches significantly contribute to the overall performance, with the MUL branch being the most crucial, which is reasonable given that image fusion inherently involves many non-linear operations. This highlights the necessity of incorporating diverse operations to enhance adaptive feature fusion. Results for MEF and MFF tasks, and visualizations of DW are provided in Appendix B.3 and B.4.
@@ 5 Conclusion
In this work, we propose a unified image fusion framework that effectively integrates both task-invariant and task-specific properties to enhance fusion performance and generalization across different fusion tasks. Future work will explore further improvements through dynamic expansion to other unseen fusion scenarios and more modalities, and integration of advanced learning techniques to develop stronger cross-modal representations, enhancing fusion quality and adaptability across diverse applications.
Acknowledgements
The research was supported by the National Natural Science Foundation of China (U23B2009, 62471158).
@@ Appendix A Supplementary Details
A.1 Additional Details of Task-specific Adaptation
For the details of three parallel operation branches in OAF module: (i) HPF branch: Conditioned by the operand $P_{h}\in\mathbb{R}^{\frac{N}{H\times W}\times K^{2}\times H\times W}$ predicted by the hyper-network $\phi_{\theta_{h}}(\cdot)$, this branch employs spatially-variant convolution on the input feature $X\in\mathbb{R}^{\frac{N}{H\times W}\times D\times H\times W}$ to obtain the resultant feature $\hat{X}_{h}\in\mathbb{R}^{\frac{N}{H\times W}\times D\times H\times W}$, where $H,W,K$ denote the height, width and the kernel size respectively. (ii) ADD branch: Conditioned by the operand $P_{a}\in\mathbb{R}^{\frac{N}{H\times W}\times D\times H\times W}$ predicted by the hyper-network $\phi_{\theta_{a}}(\cdot)$, this branch performs addition on the input feature to obtain the resultant feature $\hat{X}_{a}\in\mathbb{R}^{\frac{N}{H\times W}\times D\times H\times W}$. (iii) MUL branch: Conditioned by the operand $P_{m}\in\mathbb{R}^{\frac{N}{H\times W}\times D\times H\times W}$ predicted by the hyper-network $\phi_{\theta_{m}}(\cdot)$, this branch performs Hadamard (element-wise) multiplication on the input feature to obtain the resultant feature $\hat{X}_{m}\in\mathbb{R}^{\frac{N}{H\times W}\times D\times H\times W}$. And all the hyper-networks used are two-layer MLPs.
A.2 Optimization Objectives
@@
A.3 Additional Details of FAMO
@@ A.4 Token Exchange
where $\mathbb{I}$ is an indicator, and the threshold $\gamma$ is set to 0.020.020.02 according to the paper.
@@ Appendix B More Results and Analysis
Table 7: The ablation study for MEF task.
@@
Table 8: The ablation study for MFF task.
@@
B.1 Ablation Study for MEF and MFF Tasks
@@
Table 9: The ablation study on task-invariant integration for MEF task.
MI FMI Qabf VIF
SF 5.2338 0.8976 0.7154 1.3061
IrSF 5.2551 0.8975 0.7125 1.3059
IeSF 5.2512 0.8977 0.7163 1.3074
TE 5.2662 0.8969 0.7096 1.3155
PA 5.2428 0.8975 0.7098 1.3056
IPA 5.2273 0.8977 0.7116 1.3048
Table 10: The ablation study on task-invariant integration for MFF task.
MI FMI Qabf VIF
SF 6.2303 0.8822 0.6455 1.5997
IrSF 6.1816 0.8818 0.6415 1.5916
IeSF 6.2373 0.8822 0.6516 1.6019
TE 6.2386 0.8817 0.6467 1.5952
PA 6.2835 0.8825 0.6461 1.6088
IPA 6.2690 0.8824 0.6541 1.6091
B.2 Analysis on Task-invariant Integration for MEF and MFF tasks
We conduct ablation studies to evaluate the effectiveness of IeSF and IPA in Task-invariant Integration for MEF and MFF tasks. As shown in Tab. 9 and Tab. 10, by replacing the IeSF with IrSF or replacing IPA with TE, we observe drops in overall performance, demonstrating the necessity of both IeSF and IPA modules.
Table 11: The ablation study on task-specific adaptation for MEF task.
MI FMI Qabf VIF
W/o HPF 6.0651 0.8999 0.7247 1.4957
W/o ADD 5.8250 0.8995 0.7215 1.4760
W/o MUL 6.1305 0.8984 0.7099 1.5164
W/o DW 4.8277 0.8987 0.6894 1.2815
Ours 6.2073 0.9000 0.7227 1.5338
Table 12: The ablation study on task-specific adaptation for MFF task.
MI FMI Qabf VIF
W/o HPF 6.2898 0.8833 0.6801 1.6039
W/o ADD 6.3395 0.8845 0.6831 1.6385
W/o MUL 6.3512 0.8831 0.6663 1.6148
W/o DW 6.2616 0.8840 0.6862 1.6245
Ours 6.5463 0.8847 0.6973 1.6371
B.3 Analysis on Task-specific Adaptation for MEF and MFF tasks
We conduct ablation studies to evaluate the effectiveness of three branches in Task-adaptive Adaptation for MEF and MFF tasks. As shown in Tab. 11 and Tab. 12, while all branches contribute to the overall performance improvement, their contributions differ. For example, the MFF task relies more heavily on the HPF branch, as clarity is directly related to high-frequency details.
@@ B.3 Analysis on Task-specific Adaptation for MEF and MFF tasks
B.4 Visualization of TA
The weights for each branch of OAF for three fusion tasks is visualized in Tab. 13. It can be observed that the weight assigned to the HPF branch is relatively small, as high-frequency details comprise only a minor portion of the overall information. Nevertheless, as demonstrated in the ablation study on task-specific adaptation, these high-frequency components play a critical role in boosting performance.